OpenStudy (anonymous):

w = xe^y /z, x = t^6, y = 9 - t, z = 4 + 8t

OpenStudy (anonymous):

total differential?

OpenStudy (anonymous):

Its the chain rule to find dw/dt

OpenStudy (anonymous):

multivariable chain rule is: \[dw = f_xdx +f_ydy +f_zdz \] divide by dt to get \[\large \frac{dw}{dt}=f_x \frac{dx}{dt}+f_y\frac{dy}{dt}+f_z\frac{dz}{dt} \]

OpenStudy (anonymous):

I keep getting the wrong answer for my homework

OpenStudy (anonymous):

the rest is only substitution, \[ \frac{dx}{dt}= 6t^5 \] \[ \frac{dy}{dt}=-1\] \[ \frac{dz}{dt}=8 \]

OpenStudy (anonymous):

and the rest are partial derivatives of the function itself.